OpenStudy (anonymous):

solve by the quadratic formula and simplify answer: x^2-3x-1=0

OpenStudy (anonymous):

a=1 b=-3 c=-1 plug into: \[\frac{ -b \pm \sqrt{b ^{2}-4ac } }{ 2a }\]

jhonyy9 (jhonyy9):

first calcule the discriminant with formula D= sqrt(b^2 -4ac) so and after this x1 ,x2 = (-b +/- sqrt(b^2 -4ac))/2a = ... hope that you have understood it right good luck bye